. A student was given two metal cubes that looked similar. One was 1.05 cm on an edge and had a mass of 14.32 grams; the other was 2.66 cm on a side and had a mass of 215.3 grams. Determine whether these two cubes of metal are the same material using only the data given. 


1
Expert's answer
2021-06-22T07:13:43-0400

Density, "D= m\/v"

D1= 14.32/1.053

=14.32/1.16 =12.34g/cm3

D2= 215.3/2.663

=215.3/18.82 =11.44g/cm3

The cubes are not of the same material